OpenStudy (anonymous):

A desk is on sale for$496, which is 36% less than the regular price. What is the regular price?

OpenStudy (anonymous):

P - 0.36P = 496 solve for P

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that's not right @sourwing

OpenStudy (anonymous):

100%-36%=64% 496=0.64 496/0.64=775

OpenStudy (anonymous):

well @saraward070911 already said the answer she's correct

OpenStudy (anonymous):

@micXD13 how's it not right?

OpenStudy (anonymous):

look at @saraward070911 explanation

OpenStudy (anonymous):

My way of solving is exactly the same as hers P - 0.36P = 496 0.64P = 496 P = 775

OpenStudy (anonymous):

sorry you are correct my explanation looked like hers and i got confused by your answer

OpenStudy (anonymous):

i didn't see the solve for p part

OpenStudy (anonymous):

P is 0.64

OpenStudy (anonymous):

100%-36%=64% 496=0.64 496/0.64=775